25.0 g piece of aluminum is heated to 82.4 °C and dropped into a calorimeter containing water at 22.3 °C. The final temperature of the water and metal is 24.9 °C. Calculate the mass (in grams) of the water in the calorimeter.
1
Expert's answer
2019-09-30T05:14:10-0400

q = m*c*dT


First I converted 24.03J/°Cmol to grams

(24.03J/°Cmol) * (1 mol Al/27g Al)= 0.89J/°Cg


The final temp of water will be the same as the final temp of aluminum


q(aluminum) = 25g * 0.89J/°Cg * (24.9°C - 82.4°C) = -1279.375 J


q(water) = +1279.375 J = m * 4.18J/g°C * (24.9°C - 22.3°C)


1279.375 = 10.868m

m = 117.72 grams of water which is almost 118g.